Consider the number 35964

How many 3 digit numbers can be formed using digits from 35964 if no digits may be repeated? What is the sum on all of those 3 digit numbers?
Answer #1

total number of digits in number 35964 =5

hence  digit numbers can be formed using digits from 35964 =5*4*3 =60

as each digit appears =60*3/5=36 times

therefore  sum on all of those 3 digit numbers =36*(3+5+9+6+4)=972